OpenStudy (albert0898):

A girl can ski down a hill five times as fast as she can climb up the same hill. If she can climb up the hill and ski down in a total of 9 minutes, how many minutes does it take her to climb up the hill? Please explain in full detail, and explain any shortcuts. Thank you. --Taken from Barron's Let's Review: Integrated Algebra [1] 1.8 [2] 4.5 [3] 7.2 [4] 7.5 Answer = [4] = 7.5 How do you get 7.5?

OpenStudy (anonymous):

Define your variables and set up an equation. D = time down the hill U = time up the hill U = 5D D + U = 9 So D + 5D = 9

OpenStudy (p0sitr0n):

x =time to ski 5x = time to climb time = 9minutes x + 5x =9 6x = 9 x =1.5 1.5 *5 = climb = 7.5